Falconmaster:

Which of the following are photosynthetic autotrophs? algae plants all bacteria some bacteria arthropods some protists

Warriorz13:

Hmmmmmm, I wanna say some bacteria, algae, plants.

JustSaiyan:

Photosynthetic Autotrophs are organisms that use light to make energy. These would include Algae, Plants, and some bacteria, and some protists

Warriorz13:

Darn protists

JustSaiyan:

Yes. they are like the misfits in the cafeteria.

Warriorz13:

They are me x'D

Falconmaster:

1 +1 = jew

JustSaiyan:

1+1=2^1

Warriorz13:

BOI

Warriorz13:

1+1=WINDOW